An old fashion country school room has double desks. If each student in the class sits at a different desk, there are 2 students without seats. If 2 students sit together at a double desk, there is one empty desk. How many students and how many desks are there in the class?
Answer by ankor@dixie-net.com(22740)

An old fashion country school room has double desks.
If each student in the class sits at a different desk, there are 2 students without seats.
If 2 students sit together at a double desk, there is one empty desk.
How many students and how many desks are there in the class?
:
Let s = number of students
Let d = number of desks
:
"If each student in the class sits at a different desk, there are 2 students without seats."
therefore
s = d + 2
:
" If 2 students sit together at a double desk, there is one empty desk."
s = 2(d-1)
s = 2d - 2
Replace s with (d+2) from the 1st statement
d + 2 = 2d - 2
2 + 2 = 2d - d
d = 4 desks
and
s = 4 + 2
s = 6 students
:
I'll let you check this in the two statements and see if it is true.